Scott's family grass seed sells for $2.45 per pound and Scott's Spot Filler grass sells for $2.10 per pound. How many pounds of each should be mixed to get a 10 pound mixture that sells for approx. $2.20.
Answer by stanbon(75887)

Scott's family grass seed sells for $2.45 per pound and Scott's Spot Filler grass sells for $2.10 per pound. How many pounds of each should be mixed to get a 10 pound mixture that sells for approx. $2.20.
-----------------------
EQUATIONS:
Quantity Eq.: x + y = 10 lbs
Value Eq..: 2.45x + 2.10y = 2.20*10
------------------
Rearrange:
245x + 210y = 2200
210x + 210y = 2100
--------------------
35x = 100
x = 2.857 lb of $2.45 seed
y = 10-2.857 = 7.1428 lb of $2.10 seed
===========================
Cheers,
Stan H.